Gilmore, Inc., had equity of $145,000 at the beginning of the year. At the end of the year, the company had total assets of $210,000. During the year, the company sold no new equity. Net income for the year was $27,000 and dividends were $5,800. a. Calculate the internal growth rate for the company. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Let’s first calculate return on equity (ROA)

Return on Asset (ROA) = Net income / Total Assets   

Where,

Net income = $27,000

Total Asset at the end of period = $210,000

Dividend paid = $5,800

Retained earnings = Net income - Dividend paid = $27,000 - $5,800

= $21,200

Total Asset at the beginning of period = Total Asset at the end of period - retained earnings

= $210,000 - $21,200

= $188,800

Therefore, Average asset = (Total Asset at the end of period +Total Asset at the beginning of period)/2

= ($210,000 + $188,800)/2

= $199,400

Therefore,

Return on Asset (ROA) = $27,000 / $199,400 = 0.1354 or 13.54%

Now,

Internal Growth rate = Retention rate * ROA

Retention rate (b) = (1- payout ratio)

Where, Payout ratio = dividend paid / net income = $5,800 / $27,000 = 0.2148 or 21.48%

Therefore, retention rate = (1- payout ratio) = (1-0.2148) = 0.7852

Therefore, Internal Growth rate by using average asset

Internal Growth rate =0.7852 * 13.54% = 10.63%